Calculate the molarity of the following solution.

49.73g H2SO4 dissolved in 500.0mL of solution

(Please show all steps. Thank you.)
Answer #1

Molarity = number of moles / volume of solution in L

number of moles = 49.73 g / 98.079 g/mol = 0.507 mole

Molarity = 0.507 / 0.5 = 1.01 M

Therefore, the molarity of solution = 1.01 M